How do you solve the linear equation of x/8 +2=14 ? i really have no clue where to start.
Found 2 solutions by Alan3354, AmparoV:
Answer by Alan3354(69443)

How do you solve the linear equation of x/8 +2=14 ?
----------
x is the only literal, so I'll assume you want to solve for x.
x/8 +2=14
Subtract 2 from both sides
x/8 = 12
Multiply both sides by 9
x = 96
We're done.
No animals were harmed in this solution.

Answer by AmparoV(10)

x/8+2=14
First you would subtract two to both sides so you're left with x/8=12
Then you times both sides by 8 it would cancel on the left so you're left with x=96
Hope this helps